OpenStudy (anonymous):

sin(2t)=sqrt(2)cos(t)

OpenStudy (math&ing001):

sin(2t)=2sin(t)cos(t)=sqrt(2)cos(t) So 2sin(t)=sqrt(2) ie sin(t)=sqrt(2)/2 Should be easier now ;)

OpenStudy (anonymous):

@math&ing001 @nkb130 You cannot divide both sides by cos t. Otherwise you lose "possible solutions". You should subtract sqrt(2) cos(t) from both sides, and set equal to 0. Now factor out cos(t) and set both factors equal to 0, so you get all solutions for this equation.

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

@Easyaspi314 you're right I forgot about that. Should be cos(t)=0 or sin(t)=sqrt(2)/2 Thx for the reminder.

OpenStudy (anonymous):

@math&ing001 Welcome.